A deal has two kinds of mets; one costs 90 cents a pound, the other 60 cents a pound. He wishes to make 50 pounds of a mixture that will cost 72 cents a pound. How many pounds of each kind should he use?
Answer by ptaylor(2048)

Let x=amount of 90 cent meat needed
Then 50-x=amount of 60 cent meat needed
Now we know that the value of the meat before it's mixed ((0.90x+0.60(50-x)) has to equal the value of the meat after it's mixed (50*0.72). So our equation to solve is:
0.90x+0.60(50-x)=50*0.72 get rid of parens
0.90x+30-0.60x=36 subtract 30 from each side
0.90x+30-30-0.60x=36-30 collect like terms
0.30x=6 divide both sides by 0.30
x=20 lb---------------------------------amount of 90 cent meat needed
50-x=50-20=30 lb----------------------amount of 60 cent meat needed
CK
20*0.90+30*0.60=50*0.72
18+18=36
36=36

Hope this helps---ptaylor